Dr. Bryan Carpenter Department of Electronics and Computer Science University of Southampton Highfield Southampton S09 5NH

Tel: 0703 593081 email: dbc@uk.ac.soton.ecs

July 2, 1993

Dr. Philip Wadler, Computing Science Department, The University, Glasgow G12 8QQ

Dear Dr. Wadler,

I hope you can consider the enclosed paper, Some Lattice-based Scientific Problems, Expressed in Haskell, for publication in Journal of Functional Programming.

Yours sincerely,